OpenStudy (anonymous):

c. When the center of Earth is 2 × 1011 meters from the center of Mars, the force of gravity between the two planets is about 64.32 × 1014 newtons. The mass of Earth is about 6 × 1024 kilograms, and the mass of Mars is about 6.4 × 1023 kilograms. Using these values, estimate the gravitational constant.

OpenStudy (phi):

do you have a formula for G ?

OpenStudy (midhun.madhu1987):

The universal law of gravitation states that: \[F = GMm \div d ^{2}\] where F = Gravitational Force G= Universal Gravitational constant M = Mass of Earth m = Mass of mars d = distance between the 2 objects

OpenStudy (midhun.madhu1987):

You can find G by substituting other values..

OpenStudy (midhun.madhu1987):



OpenStudy (anonymous):

the formula I'm using for G is G=Fg d2/m*M

OpenStudy (anonymous):

@phi

OpenStudy (phi):

now put in numbers in place of the symbols

OpenStudy (phi):

2 × 1011 meters from the center of Mars, the force of gravity between the two planets is about 64.32 × 1014 newtons. The mass of Earth is about 6 × 1024 kilograms, and the mass of Mars is about 6.4 × 1023 kilograms. \[ G = \frac{F_g \cdot d^2}{m\cdot M}\\ G= \frac{64.32\cdot 10^{14} \ N \cdot (2\cdot 10^{11})^2 \ m^2}{6\cdot10^{24} \cdot 6.4\cdot 10^{23} \ kg^2} \] are you sure about the force being 64.32 x 10^14 (normally the leading number would be 6.432)?

OpenStudy (anonymous):

sorry typo, I meant 6.432

OpenStudy (phi):

so the expression is \[ G= \frac{6.432\cdot 10^{14} \ N \cdot (2\cdot 10^{11})^2 \ m^2}{6\cdot10^{24} \cdot 6.4\cdot 10^{23} \ kg^2}\]

OpenStudy (phi):

first, square 2*10^11 which is 2*10^11 * 2 * 10^11 or, reordering 2*2 * 10^11*10^11 I assume you know 2*2 is 4 what is 10^11 * 10^11 ?

OpenStudy (anonymous):

1*10^22

OpenStudy (phi):

10*10 is 100 or using exponents \( 10^1 \cdot 10^1 = 10^2 \) or 100*100= 10000 using exponents \(10^2 \cdot 10^2 = 10^4 \) do you see a pattern ? if you have the same base (and we do, it is 10) then when you multiply them, you add their exponents

OpenStudy (anonymous):

okay

OpenStudy (phi):

\[ G= \frac{6.432\cdot 10^{14} \ N \cdot 4\cdot 10^{22} \ m^2}{6\cdot10^{24} \cdot 6.4\cdot 10^{23} \ kg^2} \]

OpenStudy (phi):

Let's just do the "numbers" (the leading coefficients) what is (6.432*4)/(6*6.4) ?

OpenStudy (anonymous):

0.67

OpenStudy (phi):

ok. now let's do the top exponent part 10^14 * 10^22 = ?

OpenStudy (anonymous):

10^36

OpenStudy (phi):

now do the bottom exponent part. what do we get ?

OpenStudy (phi):

10^24 * 10^23 = ?

OpenStudy (anonymous):

10^47

OpenStudy (phi):

so we now have \[ 0.67 \cdot \frac{10^{36}}{10^{47}} \] when you divide numbers with the same base, do top exponent - bottom exponent can you do that ?

OpenStudy (anonymous):

divide the top by the bottom?

OpenStudy (phi):

yes, but to do that you do 36-47 to get the new exponent

OpenStudy (phi):

for example \[ \frac{100}{10} = 10 \] using exponents \[ \frac{10^2}{10^1} = 10^{2-1} = 10^1 = 10 \]

OpenStudy (anonymous):

0.765

OpenStudy (phi):

? do 10^36 / 10^47

OpenStudy (phi):

the answer is 10 ^ new exponent the new exponent is 36 - 47

OpenStudy (anonymous):

-11

OpenStudy (phi):

yes, \(10^{-11}\)

OpenStudy (phi):

so we now have \[ 0.67 \cdot \frac{10^{36}}{10^{47}} \\ 0.67 \cdot 10^{-11} \] now we put the answer in standard form. we want 6.7 instead of 0.67 to do that we multiply by 10. and then divide by 10, like this \[ 0.67 \cdot 10 \cdot \frac{10^{-11}}{10^1} \]

OpenStudy (phi):

0.67*10 is 6.7 we have \[ 6.7 \cdot \frac{10^{-11}}{10^1} \] can you do the 10 part? remember : top exponent - bottom exponent

OpenStudy (anonymous):

1^-12

OpenStudy (phi):

yes 10^-12 so the answer is 6.7 x 10^-12

OpenStudy (anonymous):

now I need to rewrite the formula to solve for one of the mass values

OpenStudy (phi):

if you start with \[ G = \frac{F_g \cdot d^2}{m\cdot M} \] multiply both sides by m (which means write m * on both sides) can you do that?

OpenStudy (anonymous):

how would I do that?

OpenStudy (phi):

multiplying by a letter (like "m") is easy... you write m on both sides example, say you have x=y and you multiply both sides by m. you write mx = my

OpenStudy (anonymous):

okay got it

OpenStudy (phi):

like this \[ mG = m\frac{F_g \cdot d^2}{m\cdot M}\] you can write the right-hand side as \[ mG =\frac{m}{m} \frac{F_g \cdot d^2}{ M}\]

OpenStudy (phi):

any idea what m/m is ?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

m?

OpenStudy (phi):

if it were numbers: 2/2 or 3/3 or 7/7 ?

OpenStudy (anonymous):

1

OpenStudy (phi):

yes. anything divided by itself is 1 (which is why I did what we did) so we have \[ mG =\frac{m}{m} \frac{F_g \cdot d^2}{ M} \\ mG =1\cdot\frac{F_g \cdot d^2}{ M} \\ mG =\frac{F_g \cdot d^2}{ M}\] 1 times anything is the anything, which is how we get the last line

OpenStudy (phi):

now multiply both sides by \(\frac{1}{G} \)

OpenStudy (anonymous):

mG/g=Fg*d^2/m*1/G

OpenStudy (phi):

it is clearer if you use G (not g) on the left side on the left side you have \[ \frac{mG}{G} \] what happens to the G/G ?

OpenStudy (anonymous):

it's 1

OpenStudy (phi):

and what does 1* m simplify to ?

OpenStudy (anonymous):

1m

OpenStudy (phi):

or just m

OpenStudy (phi):

so you have \[ m= \frac{F_g \cdot d^2}{ M}\cdot \frac{1}{G} \] on the right side, to multiply fractions, you multiply top*top and bottom * bottom

OpenStudy (anonymous):

okay

OpenStudy (phi):

in other words you get \[ m=\frac{F_g \cdot d^2}{G\cdot M} \]

OpenStudy (phi):

the bottom could also be written M*G. you can change the order when multiplying

OpenStudy (anonymous):

okay